### v1

@@ 3 Main Results
Our game admits many equilibria because each agent may possess arbitrarily many signals, allowing their successor’s belief to become arbitrarily pessimistic at off-path histories. This in turn discourages agents from revealing those off-path histories, making those beliefs self-fulfilling. Proposition 1 characterizes some of those equilibria:
@@
Fix $(u,v,\mathbf{f},\pi_{0})$. For every $\delta\in(0,1)$, $i\in\{1,2,...,l\}$ with $L_{i}>1$, and $n\in\mathbb{N}\cup\{\infty\}$, there is a strict equilibrium where (i) agents never disclose any $s\neq s_{i}$ and (ii) the number of $s_{i}$ disclosed at history $\mathbf{h}\equiv\{h_{1},...,h_{l}\}$ and private signal $s_{j}$ is $\min\{n,h_{i}+\mathbb{I}\{i=j\}\}$.
The proof is in Appendix A. Proposition 1 implies that regardless of $\delta$, there is always an equilibrium where agents disclose all realized $s_{1}$ and nothing else. However, it also implies that for any $s_{i}$ with a likelihood ratio more than $1$, there is also a strict equilibrium in which agents never disclose anything other than $s_{i}$ and disclose up to $n$ realizations of $s_{i}$. Such equilibria exist even if there are other signals with strictly higher likelihood ratios than $s_{i}$, as they can be sustained by the off-path belief that as long as any signal other than $s_{i}$ (including those with higher likelihood ratios) is disclosed, the agent who disclosed it must have concealed many signals with low likelihood ratios, which justifies their successors’ pessimistic beliefs about $\theta$.
We view such beliefs as unreasonable since agents should become more optimistic rather than pessimistic once their predecessors replace a signal with a lower likelihood ratio with one that has a higher likelihood ratio. We are unaware of standard refinements in signaling games that can rule out those unreasonable equilibria. Hence, we propose two refinements, each of which is sufficient to rule out those unreasonable equilibria. [^6] By the end of this section, we will discuss the relationship between our refinements and the truth-leaning refinement in Hart, Kremer, and Perry (2017) as well as the receiver-optimal equilibria in Rappoport (2025). Analogs of our refinements are also used in other dynamic disclosure models such as the one in Guttman, Kremer, and Skrzypacz (2014).
Our first refinement is on agents’ equilibrium strategy, and in particular, the component of their strategy that specifies the set of signals disclosed to the next agent upon observing each history $\mathbf{h}$ and private signal $s$, which we denote by $\sigma_{D}(\mathbf{h},s)\in H$. For every vector $\mathbf{h}\in H$ and $j\in\{0,1,...,l\}$, let $\mathbf{h}[j]\in H$ denote an $l$-dimensional vector where the first $j$ entries coincide with $\mathbf{h}$ and the other entries are all 0. By definition, $\mathbf{h}[0]=\mathbf{0}$ and $\mathbf{h}[l]=\mathbf{h}$ for every $\mathbf{h}\in H$.
  Threshold Strategy.
We view our threshold strategy as a natural extension of the equilibrium strategies in Dye (1985) to our dynamic setting. Intuitively, a threshold strategy is one where at every history-signal pair, agents reveal all available signals (including his private signal and the signals contained in his history) with likelihood ratios above some cutoff and conceal all other signals. It allows the disclosure threshold to vary across histories and private signals but rules out strategies where at some history-signal pair, agents disclose $s_{2}$ but conceal $s_{1}$. For a concrete example, suppose $J(\mathbf{0},s)=1$ for every $s\in S$ and $J(\mathbf{h},s)=0$ for every $\mathbf{h}\neq\mathbf{0}$ and $s\in S$, then agents disclose $s_{1}$ when the length of his history is 0 and disclose nothing otherwise.
Our next refinement is on the belief map $\pi:H\rightarrow[0,1]$, which stands for the probability agents assign to state $\overline{\theta}$ after observing their history but before observing their private signal.
@@
Intuitively, a belief map is monotone if fixing the total number of signals disclosed, agent $t+1$ will become more optimistic about the state once agent $t$ replaces a signal that has a lower likelihood ratio with one that has a higher likelihood ratio. This property is satisfied in the equilibria of Dye (1985)’s model where the sender has at most one signal.
Among the equilibria described in Proposition 1, the one where agents reveal all realized $s_{1}$ but nothing else satisfies both of our refinements. The ones where agents disclose $s_{i}$ for some $i\geq 2$ but conceal $s_{1}$ fail both refinements. However, it remains unclear whether there are equilibria where agents reveal $\{s_{1},s_{2},...,s_{i}\}$ for some $i\geq 2$ and conceal other signals, or whether there are equilibria with a monotone belief map where agents reveal signals other than $s_{1}$.
Our main result, Theorem 1, shows that when communication frictions are small enough, none of the above can happen in any strict equilibrium that satisfies at least one of our refinements.
@@
    Signals $s_{2},...,s_{l}$ are never revealed in any strict equilibrium with a threshold strategy. [^7] In any equilibrium with a threshold strategy, agents will use a threshold strategy in equilibrium but need to have incentives not to deviate to any other strategy, including strategies that are not threshold strategies.
@@
    Signals $s_{2},...,s_{l}$ are never revealed in any strict equilibrium with a monotone belief map.
@@
    There are only two strict equilibria in threshold strategies: one where agents never disclose anything and one where agents only disclose signal $s_{1}$ and disclose all realized $s_{1}$.
@@
    Every strict equilibrium with a monotone belief map is characterized by an integer $n\in\mathbb{N}\cup\{\infty\}$ such that $\sigma_{D}(\mathbf{h},s_{i})=\min\{n\mathbf{1}_{1},(\mathbf{h}+\mathbf{1}_{i})[1]\}$ for every $\mathbf{h}\in H$ and $i\in\{1,2,...,l\}$, i.e., agents conceal all signals other than $s_{1}$ and disclose up to $n$ realizations of $s_{1}$.
The proof of Theorem 1 and Corollary 1 are both in Section 4.
In order to better visualize how communication frictions affect the selectiveness of agents’ disclosure, our next result allows for any value of $\delta$ and provides a full characterization of equilibria in which agents use constant threshold strategies. These are strategies where the disclosure threshold is independent of agents’ histories and private signals.
  Constant Threshold Strategy.
A strategy $\sigma$ is a constant threshold strategy if there exists $j\in\{0,1,...,l\}$ such that $\sigma_{D}(\mathbf{h},s_{i})=(\mathbf{h}+\mathbf{1}_{i})[j]$ for every $\mathbf{h}\in H$ and $i\in\{1,2,...,l\}$.
For example, agents will never disclose any signal when the threshold is constantly 0, and will disclose all signals when the threshold is constantly $l$. Theorem 2 focuses on equilibria in constant threshold strategies and shows that as $\delta$ increases, agents become increasingly selective in disclosing information, in the sense that a smaller set of signals can be disclosed in equilibrium.
@@
    If $L_{i}>1$, then there exists $\delta_{i}\in(0,1)$ such that there is an equilibrium where agents use a constant threshold $i$ if and only if $\delta\leq\delta_{i}$. Furthermore, $\delta_{i}$ is strictly decreasing in $i$.
@@
    If $L_{i}\leq 1$, then agents will never use a constant threshold $i$ regardless of $\delta$.
Theorem 2 implies that (i) signals with likelihood ratio no more than $1$ will never be disclosed in any equilibrium with a constant threshold strategy and (ii) any signal with likelihood ratio more than $1$ will be disclosed in some equilibria with constant threshold strategies if and only if the communication success rate $\delta$ is below some cutoff. Furthermore, the cutoff $\delta$ is strictly increasing in the likelihood ratio of the signal. Since constant threshold strategies are threshold strategies, we know that when $\delta>\delta_{2}$, the only equilibrium where agents use a constant threshold strategy is the one where agents disclose a signal if and only if it is $s_{1}$.
Our finding that disclosure becomes more selective as $\delta$ increases stands in contrast to the results in Dye (1985). In Dye (1985)’s model, as the probability that the sender has evidence increases, he becomes less selective in the sense that a larger set of realized signals are disclosed in equilibrium. In the limiting case studied by Grossman (1981) and Milgrom (1981) where the sender has evidence for sure, no disclosure leads to the receiver to infer that the sender has the worst possible signal, which motivates the sender to disclose even the second worst signal.
In our model, a larger $\delta$ increases the chances that agents observe the signals passed on to them by their predecessors, so absent strategic concerns, their successors should expect them to have more signals. However, disclosure becomes more selective due to the nature of the uncertainty their successors face about their information structures, which differs from that in Dye (1985).
To understand the differences, suppose agents are expected to disclose signals in $\{s_{1},...,s_{i}\}$ (i.e., good signals) and to conceal signals in $\{s_{i+1},...,s_{l}\}$ (i.e., bad signals). After the last communication failure, the sequence of realized signals consists of a number of good signals as well as some subsequences of bad signals (i.e., bad subsequences), where the length of each bad subsequence can be anything from 0 to $\infty$. Each agent observes the number of good signals after the latest communication failure but not the lengths of the bad subsequences. When he observes $k$ good signals, he expects there to be $k+1$ bad subsequences: one before the first good signal, one after the last good signal, and one in between each pair of adjacent good signals. An example of such a sequence of signals is depicted in Figure 1, which helps to understand why the number of bad subsequences equals the number of good signals plus $1$.
@@
Figure 1: A sequence of realized private signals after the most recent communication failure, with red circles representing bad signals and green circles representing good signals. In this example, there are 44 good signals and $5$ bad subsequences. The bad subsequence before good signal $1$ has length 33, the bad sequence between good signals $1$ and 22 has length 0, the bad subsequence between good signals $2$ and 33 has length 22, the bad subsequence between good signals $3$ and 44 has length 66, and the bad subsequence after good signal $4$ has length 22.
An important observation is that, conditional on the state, agents’ belief about the length of each bad subsequence is independent of the number of good signals. Therefore, upon disclosing another good signal, there are two effects on the next agent’s belief. There is a direct effect that is positive whenever the likelihood ratio of that good signal exceeds $1$, but there is also an indirect effect since revealing another good signal also implies that there is an additional bad subsequence, and bad subsequences are more likely to arise when the state is $\underline{\theta}$. This indirect effect is absent in Dye (1985)’s model where disclosing a signal shows to the receiver that no information is hidden.
Therefore, the agents in our model have incentives to disclose a signal if and only if its likelihood ratio is large enough so that the direct effect dominates the indirect effect. As $\delta$ increases, the expected length of each bad subsequence increases conditional on each state, so the indirect effect of an additional bad subsequence becomes stronger yet the direct effect remains unchanged. Hence, agents become more selective in disclosing signals as fewer signals will have high enough likelihood ratios to overcome the increasingly pronounced indirect effect.
In what follows, we present a proof of Theorem 2, which helps readers to visualize the two effects of disclosing an additional signal, as well as how they depend on $\delta$.
@@
Since $v(\theta,a^{t+1})$ is strictly increasing in $a^{t+1}$ and agent $t+1$’s action is strictly increasing in the probability with which his belief assigns to $\overline{\theta}$, we know that at every $(\mathbf{h},s_{i})\in H\times S$, agent $t$ will choose a vector $\mathbf{h^{\prime}}\leq\mathbf{h}+\mathbf{1}_{i}$ in order to maximize $\pi(\mathbf{h^{\prime}})$.
Suppose all agents use a constant threshold $n$, then under the steady state history distribution in state $\theta$, the probability that an agent observes history $\mathbf{h}\equiv(h_{1},h_{2},...,h_{n},0,...,0)$ is
@@
Equation (3.5) clearly shows the two effects of disclosing an additional signal. As we mentioned earlier, when the total number of signals disclosed is $\overline{h}$, there are $\overline{h}+1$ bad subsequences (some of them may have length zero) after the most recent communication failure, which explains why the power of the second term on the right-hand-side is $\overline{h}+1$. By disclosing an additional signal $s_{i}$, the direct effect multiplies the likelihood ratio by $L_{i}$ (the likelihood ratio of $s_{i}$) and the indirect effect multiplies the likelihood ratio by
@@
which is the contribution of an additional bad subsequence. Crucially, as long as $\delta>0$ and $n<l$, the indirect effect, captured by (3.6), is always strictly less than $1$. This is because each bad subsequence only consists of signals with likelihood ratios below some cutoff. The value of (3.6) decreases in $\delta$ since a larger $\delta$ increases the expected length of each bad subsequence conditional on each state. Therefore, the value of (3.6) is bounded below by its limit where $\delta\rightarrow 1$, which equals
$\frac{\underline{f_{1}}+...+\underline{f_{n}}}{\overline{f_{1}}+...+\overline{f_{n}}}$ (3.7)
This lower bound is more than the inverse of $L_{1}$ but is less than the inverse of $L_{n}$, i.e., disclosing the signal with the highest likelihood ratio is sufficient to overcome the negative indirect effect but as long as signals other than $s_{1}$ are disclosed, disclosing the worst one among those signals is never sufficient to overcome to negative indirect effect when $\delta$ is large enough.
@@
with strict inequality for every $n<l$, we know that $L_{n}>1$ is necessary for (3.8). Since the left-hand-side of (3.8) is strictly decreasing in $\delta$, which converges to $L_{n}$ as $\delta\rightarrow 0$ and converges to (3.7) times $L_{n}$ as $\delta\rightarrow 1$, we know that for every $n\geq 2$ that satisfies $L_{n}>1$, there exists $\delta_{n}\in(0,1)$ such that inequality (3.8) holds if and only if $\delta\leq\delta_{n}$. Since $\delta_{n}$ is pinned down by the equation
@@
Theorem 2 implies that once $\delta$ exceeds the cutoff $\delta_{n}$ defined in equation (3.9), there is no equilibrium in constant threshold strategies where agents use threshold $n$ across all histories and private signals. This seems to suggest that once we focus on equilibria in constant threshold strategies, the average amount of information an agent receives is discontinuous in $\delta$ in the sense that the average amount of information drops at the cutoff values of $\delta$.
Proposition 2 shows that there is no such discontinuity. Focusing on the most informative equilibrium in constant threshold strategies, i.e., the one with the highest disclosure threshold in terms of the signal index, the average amount of information an agent receives, measured by difference in Kullback–Leibler divergence between the agent’s prior and his belief after observing his history (but before observing his private signal), is continuous and strictly increasing in $\delta$.
Formally, let $\mathbf{h}_{\delta}^{(n)}\in H$ denote the disclosed history observed by an agent in the steady state when all agents use constant threshold strategy $n$, which is a random variable. For any $\delta\in(\delta_{n+1},\delta_{n}]$, let $\mathbf{h}_{\delta}\equiv\mathbf{h}_{\delta}^{(n)}$ , so that $\mathbf{h}_{\delta}\in H$ is the disclosed history observed by an agent in the steady state when we select the equilibrium with the largest threshold $n$ that exists for each $\delta$. Theorem 2 implies $\mathbf{h}_{\delta}$ is well-defined. We denote the difference in the expected Kullback–Leibler divergence between an agent’s prior belief about the state and his posterior belief by
@@
where $g(p)\equiv-p\log p-(1-p)\log(1-p)$ is the entropy function. We refer to $I(\theta;\mathbf{h}_{\theta})$ as mutual information. Intuitively, $I(\theta;\mathbf{h}_{\theta})$ is greater when agents learn more about the state from their histories. Proposition 2 shows that the mutual information is continuous and strictly increasing in $\delta$, even when the number of signals disclosed changes at the cutoff values of $\delta$.

### v2

@@ 3 Main Results
Our game admits many equilibria because each agent may possess arbitrarily many signals, allowing their successor’s belief to become arbitrarily pessimistic at off-path histories. This in turn discourages agents from revealing those off-path histories, making those beliefs self-fulfilling. We start with some preliminary observations which describe some of those equilibria:
@@
Fix $(u,v,\mathbf{f},\pi_{0})$. For every $\delta\in(0,1)$, $i\in\{1,2,...,l\}$ with $L_{i}>1$, and $n\in\mathbb{N}\cup\{\infty\}$, there is a strict equilibrium where (i) agents never disclose any $s\neq s_{i}$ and (ii) the number of $s_{i}$ disclosed at history $\mathbf{h}=(h_{1},...,h_{l})$ and private signal $s_{j}$ is $\min\{n,h_{i}+\mathds{1}\{i=j\}\}$.
The proof is in Appendix A.2. Proposition 1 implies that regardless of $\delta$, there is always an equilibrium where agents disclose all realized $s_{1}$ and nothing else. However, it also implies that for any $s_{i}$ with a likelihood ratio more than $1$, there is also a strict equilibrium in which agents never disclose anything other than $s_{i}$ and disclose up to $n$ realizations of $s_{i}$. Such equilibria exist even if there are other signals with strictly higher likelihood ratios than $s_{i}$, as they can be sustained by the off-path belief that as long as any signal other than $s_{i}$ (including those with higher likelihood ratios) is disclosed, the agent who disclosed it must have concealed many signals with low likelihood ratios, which justifies their successors’ pessimistic beliefs about $\theta$.
Such beliefs seem unreasonable since agents should become more optimistic about the state once they inherit one more high-likelihood-ratio signal and one fewer low-likelihood-ratio signal. We are unaware of standard refinements in signaling games that can rule out those unreasonable equilibria. Hence, we propose two refinements, each of which is sufficient to rule them out. [^9] Analogs of our refinements are also used in Guttman, Kremer, and Skrzypacz (2014). By the end of this section, we will discuss the relationship between our refinements and the truth-leaning refinement in Hart, Kremer, and Perry (2017) as well as the receiver-optimal equilibria in Rappoport (2025) and Gao (2025).
Our first refinement is on the belief map $\pi:H\rightarrow[0,1]$, which stands for the probability agents assign to state $\overline{\theta}$ after observing their history but before observing their private signal.
@@
An equilibrium is a monotone equilibrium if it has a monotone belief map. Intuitively, a belief map is monotone if fixing the total number of signals disclosed by agent $t$, agent $t+1$ will become more optimistic about the state once agent $t$ replaces a signal that has a lower likelihood ratio with one that has a higher likelihood ratio. Our monotonicity property is satisfied in all equilibria of Dye (1985)’s model where the sender either has one signal or has no signal.
Our second refinement is on agents’ equilibrium disclosure behaviors. Let $\sigma_{D}(\mathbf{h},s)\in H\equiv\mathbb{N}^{l}$ denote the set of signals disclosed to the next agent upon observing history $\mathbf{h}$ and private signal $s$, which is a component of agents’ strategy $\sigma$. For every $\mathbf{h}\in H$ and $j\in\{0,1,...,l\}$, let $\mathbf{h}[j]\in H$ denote an $l$-dimensional vector where the first $j$ entries coincide with $\mathbf{h}$ and the other entries are all 0. According to our definition, $\mathbf{h}[0]=\mathbf{0}$ and $\mathbf{h}[l]=\mathbf{h}$ for every $\mathbf{h}\in H$.
  Threshold Strategy.
An equilibrium is a threshold equilibrium if agents use a threshold strategy. [^10] In any threshold equilibrium, agents will use a threshold strategy in equilibrium but need to have incentives not to deviate to any other strategy, including strategies that are not threshold strategies. We view our threshold strategy as a natural extension of the equilibrium strategies in Dye (1985) to our dynamic setting. Intuitively, a threshold strategy is one where at every history-signal pair, agents reveal all available signals (including his private signal and the signals contained in his history) with likelihood ratios above some cutoff and conceal all signals with likelihood ratio below the cutoff. It allows the disclosure threshold to vary across histories and private signals. For example, suppose $J(\mathbf{0},s)=1$ for every $s\in S$ and $J(\mathbf{h},s)=0$ for every $\mathbf{h}\neq\mathbf{0}$ and $s\in S$, then agents disclose $s_{1}$ when the length of his history is 0 and disclose nothing otherwise. It rules out, for example, strategies where at some history-signal pair, agents disclose $s_{2}$ but conceal $s_{1}$.
Among the equilibria in Proposition 1, the one where agents reveal all realized $s_{1}$ but nothing else satisfies both of our refinements. The ones where agents disclose $s_{i}$ for some $i\geq 2$ but conceal $s_{1}$ fail both refinements. Some natural questions include (i) whether there are threshold equilibria where agents reveal $\{s_{1},s_{2},...,s_{i}\}$ for some $i\geq 2$ and conceal other signals, (ii) whether there are equilibria with monotone belief maps where agents reveal signals other than $s_{1}$, and (iii) how does the sender’s disclosure behavior depends on communication frictions.
Our main result, Theorem 1, shows that when communication frictions vanish, signals other than $s_{1}$ are never disclosed in any strict equilibrium that satisfies at least one of our refinements.
@@
    In any strict monotone equilibrium, signals $s_{2},...,s_{l}$ are never disclosed at any history that occurs with positive probability.
@@
    In any strict threshold equilibrium, signals $s_{2},...,s_{l}$ are never disclosed at any history that occurs with positive probability.
@@
    Each strict monotone equilibrium is characterized by some $n\in\mathbb{N}\cup\{\infty\}$ such that $\sigma_{D}(\mathbf{h},s_{i})=\min\{n\mathbf{1}_{1},(\mathbf{h}+\mathbf{1}_{i})[1]\}$ for every $\mathbf{h}\in H$ and $i\in\{1,2,...,l\}$.
@@
    For every $n\in\mathbb{N}\cup\{\infty\}$, there exists a strict monotone equilibrium where agents conceal all signals other than $s_{1}$ and disclose up to $n$ realizations of $s_{1}$.
  3. 3.
    There are only two strict threshold equilibria: one where agents never disclose anything and one where agents only disclose signal $s_{1}$ and disclose all realized $s_{1}$.
The proofs of Theorem 1 and Corollary 1 are both in Section 4.
In order to better visualize how communication frictions affect the selectiveness of agents’ disclosure, our next result allows for any value of $\delta$ and provides a full characterization of equilibria in which agents use constant-threshold strategies (i.e., constant-threshold equilibria). These are strategies where the disclosure threshold is independent of histories and private signals.
  Constant-Threshold Strategy.
A strategy $\sigma$ is a constant-threshold strategy if there exists $j\in\{0,1,...,l\}$ such that $\sigma_{D}(\mathbf{h},s_{i})=(\mathbf{h}+\mathbf{1}_{i})[j]$ for every $\mathbf{h}\in H$ and $i\in\{1,2,...,l\}$.
For example, agents will never disclose any signal when the threshold is constantly 0, and will disclose all signals when the threshold is constantly $l$. Theorem 2 focuses on constant-threshold equilibria and shows that as $\delta$ increases, agents become increasingly selective in disclosing information, in the sense that a smaller set of signals can be disclosed in this class of equilibria.
@@
    If $L_{n}>1$, then there exists $\delta_{n}\in(0,1)$ such that there is a constant-threshold equilibrium where agents use a constant threshold $n$ if and only if $\delta\leq\delta_{n}$. Furthermore, $\delta_{n}$ is strictly decreasing in $n$.
@@
    If $L_{n}\leq 1$, then for any $\delta$, agents never disclose $s_{n}$ in any constant-threshold equilibrium.
Theorem 2 implies that (i) signals with likelihood ratio no more than $1$ will never be disclosed in any constant-threshold equilibrium and (ii) a signal with likelihood ratio strictly more than $1$ will be disclosed in some constant-threshold equilibria if and only if the communication success rate $\delta$ is below some cutoff. Furthermore, $\delta_{2}>\delta_{3}>$..., which means that the cutoff $\delta$ is higher for signals with higher likelihood ratios, so higher likelihood ratio signals are disclosed in a larger range of $\delta$. Since any constant-threshold equilibrium is also a threshold equilibrium, Theorem 2 also implies that when $\delta>\delta_{2}$, there are only two constant-threshold equilibria: the one where agents never disclose anything and the one where agents disclose a signal if and only if it is $s_{1}$.
Our finding that disclosure becomes more selective as $\delta$ increases contrasts with the results in Dye (1985). In Dye (1985)’s model, as the probability that the sender has evidence increases, he becomes less selective in the sense that a larger set of realized signals are disclosed in equilibrium. In the limiting case studied by Grossman (1981) and Milgrom (1981) where the sender has evidence for sure, no disclosure leads to the receiver to infer that the sender has the worst possible signal, which motivates the sender to disclose even the second worst signal.
In our model, a larger $\delta$ increases the chances that agents observe the signals disclosed by their predecessors, so absent strategic concerns, their successors should expect them to have more signals. However, disclosure becomes more selective due to the nature of the uncertainty their successors face about their information structures, which differs from that in Dye (1985).
To understand the differences, suppose agents are expected to disclose signals in $\{s_{1},...,s_{i}\}$ (i.e., good signals) and to conceal signals in $\{s_{i+1},...,s_{l}\}$ (i.e., bad signals). After the last communication failure, the sequence of realized signals consists of some good signals and some subsequences of bad signals (i.e., bad subsequences), where the length of each bad subsequence can be anything from 0 to $\infty$ and is independent of the number of good signals in the sequence. One useful observation is that the number of bad subsequences equals the number of good signals plus $1$. An example of such a sequence of signals is depicted in Figure 1.
@@
Figure 1: A sequence of realized private signals after the most recent communication failure, with red circles representing bad signals and green circles representing good signals. In this example, there are 44 good signals and $5$ bad subsequences. The bad subsequence before good signal $1$ has length 33, the bad sequence between good signals $1$ and $2$ has length 0, the bad subsequence between good signals $2$ and $3$ has length 22, the bad subsequence between good signals $3$ and $4$ has length 66, and the bad subsequence after good signal $4$ has length 22.
When agents observe the number of good signals but not the length of each bad subsequence, disclosing another good signal shows to the next agent that there is one more good signal and one more bad subsequence. The former has a positive effect on the next agent’s belief as long as the likelihood ratio of the good signal exceeds $1$. The latter has a negative effect since bad subsequences are likely to be longer when the state is $\underline{\theta}$. This negative effect is absent in Dye (1985)’s model where disclosing any signal shows to the receiver that no information is hidden.
Therefore, the agents in our model have incentives to disclose a signal if and only if its likelihood ratio is large enough so that the direct effect dominates the indirect effect. As $\delta$ increases, the expected length of each bad subsequence increases conditional on each state, so the indirect effect of an additional bad subsequence becomes stronger yet the direct effect remains unchanged. Hence, agents become more selective in disclosing signals as fewer signals have high enough likelihood ratios to offset the increasingly pronounced indirect effect.
Next, we present a proof of Theorem 2, which helps readers to visualize mathematically the two effects of disclosing an additional signal, as well as how they depend on $\delta$.
@@
Since $v(\theta,a^{t+1})$ is strictly increasing in $a^{t+1}$ and agent $t+1$’s action is strictly increasing in the probability with which his belief assigns to $\overline{\theta}$, we know that at every $(\mathbf{h},s_{i})\in H\times S$, agent $t$ will choose a vector $\mathbf{h^{\prime}}\leq\mathbf{h}+\mathbf{1}_{i}$ that maximizes $\pi(\mathbf{h^{\prime}})$.
Suppose in equilibrium, all agents use a constant threshold strategy that discloses signals $s_{1},...,s_{n}$. Under the steady state history distribution in state $\theta$, the probability that an agent observes history $\mathbf{h}\equiv(h_{1},h_{2},...,h_{n},0,...,0)$ is
@@
Equation (3.5) clearly shows the two effects of disclosing an additional signal. As we mentioned earlier, after the most recent communication failure, when the total number of signals disclosed is $\overline{h}$, there are $\overline{h}+1$ bad subsequences (some of them may have length zero), which explains why the power of the second term on the right-hand-side is $\overline{h}+1$. By disclosing an additional signal $s_{i}$, the direct effect multiplies the likelihood ratio by $L_{i}$, as shown by the third term on the right-hand-side, and the indirect effect multiplies the likelihood ratio by
@@
which is the contribution of an additional bad subsequence. Crucially, as long as $\delta>0$ and $0<n<l$, the indirect effect, captured by (3.6), is always strictly less than $1$. This is because each bad subsequence only consists of signals with likelihood ratios below some cutoff. The value of (3.6) decreases in $\delta$ since a larger $\delta$ increases the expected length of each bad subsequence conditional on the state. Therefore, the value of (3.6) is bounded below by its limit where $\delta\rightarrow 1$, which equals
$\frac{\underline{f_{1}}+...+\underline{f_{n}}}{\overline{f_{1}}+...+\overline{f_{n}}}.$ (3.7)
This lower bound is more than the inverse of $L_{1}$ but is less than the inverse of $L_{n}$. Hence, disclosing signal $s_{1}$ (the one with the highest likelihood ratio) is sufficient to offset the indirect effect. However, as long as signals other than $s_{1}$ are disclosed, disclosing the worst one among those signals is not sufficient to overcome the negative indirect effect when $\delta$ is large enough.
@@
with strict inequality for every $0<n<l$, we know that $L_{n}>1$ is necessary for (3.8). Since the left-hand-side of (3.8) is strictly decreasing in $\delta$, which converges to $L_{n}$ as $\delta\rightarrow 0$ and converges to (3.7) times $L_{n}$ as $\delta\rightarrow 1$, we know that for every $n\geq 2$ that satisfies $L_{n}>1$, there exists $\delta_{n}\in(0,1)$ such that inequality (3.8) holds if and only if $\delta\leq\delta_{n}$. Since $\delta_{n}$ is pinned down by the equation
@@
Theorem 2 implies that once $\delta$ exceeds the cutoff $\delta_{n}$ defined in equation (3.9), there is no constant-threshold equilibrium where agents disclose signal $s_{n}$. This seems to suggest that once we focus on constant-threshold equilibria, the expected amount of information an agent receives in the steady state drops discontinuously at $\delta_{n}$.
Proposition 2 shows that there is no such discontinuity. Focusing on the most informative constant-threshold equilibrium, i.e., the one with the highest disclosure threshold in terms of the signal index, the expected amount of information an agent receives, measured by the expected difference in Kullback–Leibler divergence between the agent’s prior and his belief after observing his history (but before observing his private signal), is continuous and strictly increasing in $\delta$.
Formally, let $\mathbf{h}_{\delta}^{(n)}\in H$ denote the disclosed history observed by an agent in the steady state when all agents use constant threshold strategy $n$. For any $\delta\in(\delta_{n+1},\delta_{n}]$, let $\mathbf{h}_{\delta}\equiv\mathbf{h}_{\delta}^{(n)}$ , so that $\mathbf{h}_{\delta}\in H$ is the disclosed history observed by an agent in the steady state when we select the equilibrium with the largest threshold $n$ that exists for each $\delta$. Theorem 2 implies that $\mathbf{h}_{\delta}$ is well-defined. We denote the difference in the expected Kullback–Leibler divergence between an agent’s prior belief about the state and his posterior belief by
@@
where $g(p)\equiv-p\log p-(1-p)\log(1-p)$ is the entropy function. We refer to $I(\theta;\mathbf{h}_{\delta})$ as mutual information. Intuitively, $I(\theta;\mathbf{h}_{\delta})$ is greater when agents learn more about the state from their histories. Proposition 2 shows that the mutual information is continuous and strictly increasing in $\delta$, even when the number of signals disclosed changes at the cutoff values of $\delta$.